OpenStudy (anonymous):

A six-member committee is formed to plan a Christmas formal. The members of the group are to be selected from five teachers and nine students. If the members are randomly selected, what is the probability that it will include at least two teachers?

OpenStudy (anonymous):

the whole 6 member committee throws me off, i'm kind of rusty at stats but i'll try to help you as much as i can probability of selecting one teacher at random: 5 / (5+9) = 5 / 14 then the probability of selecting another teacher at random is: 4 / (4+9) = 4 / 13 then multiply the chance of those two events happening (5 / 14) * (4 / 13) = (20 / 182) = 10/91 = .1098 or 10.98% **that's not correct, but maybe it can spark something and help you find the right path

OpenStudy (anonymous):

***it might be correct, it might not be, sorry